Suppose a firm worth $13 million is financed with $8 million worth of debt. If the expected rate of return of the bond and the equity is 4.5% and 10.4% respectively, what's the weighted average cost of capital?_____________

Suppose your project is worth $822 with a probability of 0.9 and $492 with a probability of 1-0.9. The appropriate cost of capital is 9.2% for the overall project. If you want to raise $664 today and promise to bond investors a rate of return of $5.4, what's the expected rate of return of this bond?____________________ Please answer both Carry out calculations to at least 4 decimal places. Enter percentages as whole numbers. Example: 3.03% should be entered as 3.03. Do not include commas or dollar signs in numerical answers.
Answer #1

1)

value of debt=8 million

value of equity=13-8=5 million

what's the weighted average cost of capital=weight of debt*expected rate of return of the bond+weight of equity*cost of equity

=(8/13)*4.5%+(5/13)*10.4%

=6.7692% or 6.77%